Data concerning Kodiak National Corporation's single product appear below:

Selling price per unit $ 160.00
Variable expense per unit $ 65.60
Fixed expense per month

$ 387,040

The unit sales to attain the company's monthly target profit of $17,000 is closest to: (Round your intermediate calculations to 2 decimal places.)

6,159

2,525

4,280

4,321
Answer #1

Contribution margin=Sales-Variable cost  

=(160-65.6)=$94.4 per unit

Target Contribution margin=Fixed cost+Target profits

=387,040+17000

=$404040

Hence unit sales required=Target Contribution margin/Contribution margin per unit

=404040/94.4

=4280 units(Approx)